# XLB: A Differentiable Massively Parallel Lattice Boltzmann Library in Python

### v2

@@ 3.6 LBM Step
Moreover, leveraging the power of JAX, users can conveniently compute the gradient of a single LBM step with respect to its inputs. This is achieved using the jax.grad function, which automatically differentiates through the operations of the LBM step. By simply applying jax.grad to the step method, the gradients required for these analyses can be efficiently obtained. This feature of JAX is particularly useful for sensitivity analysis, optimization tasks, and machine learning applications.
3.7 Distributed Checkpointing, Mixed Precision, and I/O Capabilities
One of the features of the XLB library is its support for mixed precision computation, which allows users to specify different precision levels for computation and storage. For instance, a setting like f64/f32 indicates that the computation is performed in double precision (f64) while storage is handled in single precision (f32). This approach offers a balance between computational accuracy and efficiency. Lower precision storage reduces the memory footprint, enabling the simulation of larger problems and potentially enhancing performance, especially on modern GPUs.
In addition to mixed precision computation, the XLB library incorporates advanced checkpointing capabilities. Checkpointing is crucial for long-running simulations, as it allows for saving the state of a simulation at specified intervals. This feature is particularly useful for recovering from potential system failures or for pausing and resuming simulations as needed. The XLB library’s checkpointing system is designed to be distributed, meaning it can efficiently handle checkpointing in simulations that run across multiple devices or nodes.
The XLB library supports outputs in both Binary and ASCII VTK formats via the PyVista library sullivan2019pyvista and has been integrated with PhantomGaze PhantomGaze for advanced, GPU-accelerated in-situ visualizations. This integration is further enhanced by DLPack, facilitating tensor data interchange between XLB and PhantomGaze while the data remains on the GPU, circumventing the need for conventional I/O processes.
@@ 4.1 Taylor-Green Vortex in 2D
$u_{th}(\mathbf{x},t)$ $=$ $U_{0}\sin(k_{x}x)\cos(k_{y}y)\exp\left\{-\frac{t}{t_{d}}\right\}$
$v_{th}(\mathbf{x},t)$ $=$ $-U_{0}\cos(k_{x}x)\sin(k_{y}y)\exp\left\{-\frac{t}{t_{d}}\right\}$
$p_{th}(\mathbf{x},t)$ $=$ $1-\frac{\rho_{0}U_{0}^{2}}{4}\left[\cos(2k_{x}x)+\cos(2k_{y}y)\right]\exp\left\{-2\frac{t}{t_{d}}\right\}$ (9)
in which $k_{x}=2\pi/n_{x}$ and $k_{y}=2\pi/n_{y}$ with $n_{x}$ and $n_{y}$ indicating the number of cells in $x$ and $y$ respectively. Also, $t_{d}$ denotes a diffusive timescale defined as $t_{d}=\left[2\nu(k_{x}^{2}+k_{y}^{2})\right]^{-1}$.
@@
Figure 5 illustrates $\epsilon_{u}$ (in panels a, c) and $\epsilon_{\rho}$ (in panels b, d) for a range of resolutions $N=\{32,\ldots,1024\}$ using either BGK (in panels a, b) or the KBC (in panels c, d) model as the collision operator. The reported simulation results were conducted using various mixed precision ‘computation/storage’ pairs where f64 and f32 indicate double and single precision, respectively. For all cases, we used the D2Q9 lattice and periodic boundary conditions (recall that by default if no boundary condition is imposed, the streaming operation in XLB imposes periodic boundary conditions by construction). All the numerical results were obtained at $t_{f}=0.05t_{d}$ using a flow with $Re=1/\nu=1600$. We set $\Delta x=1/N$ and $\Delta t/\Delta x=0.04$ at $N=32$ where $\Delta t$ was decreased proportionally as $N$ was increased. The distribution functions $f_{i}$ were initialized based on the method of Mei:2006:ic using $u_{th}(x,t=0)$ as per (9).
The expected second order of accuracy (i.e. $\epsilon_{u}\propto N^{-2}$ and $\epsilon_{\rho}\propto N^{-2}$) is obtained for both collision models using f64/f64 precision pair. It is worth mentioning that $\epsilon_{u}$ shows a more robust convergence based on the KBC model up to $N=1024$ while the convergence of $\epsilon_{u}$ degrades for the BGK model beyond $N=256$ for f64/f64. We may attribute this to the chosen time step size or the number of iterations used for the initialization process. If the storage precision is reduced to ‘f32’ (i.e. f64/f32 cases), the second order accuracy is preserved up to a lower resolution beyond which both $\epsilon_{u}$ and $\epsilon_{\rho}$ deteriorate as $N$ increases. This is due to the fact that the results become cluttered by numerical rounding errors and hence are not improved with increasing the resolution. This exact deterioration in accuracy occurs more vividly for even lower resolutions as the computational precision is also reduced to f32 (i.e. f32/f32 cases) leading to erroneous results for higher resolutions. Consequently, LBM results with f32/f32 are not reliable unless the distribution functions $f_{i}$ values are renormalized carefully to be centered around zero lehmann2022accuracy (this feature is not currently available in XLB).
@@ 4.2 Lid-driven Cavity in 3D
Here we investigate the 3D flow inside a lid-driven cavity to further verify XLB especially under steady-state, unsteady and turbulent scenarios with non-periodic boundaries. Namely, Dirichlet boundary conditions are employed to assign no-slip condition on all side walls except for the top lid which moves horizontally with a prescribed velocity of $u(x,y,z=L/2)=(U,0,0)$. We consider a box of size $L^{3}$ containing an incompressible Newtonian fluid with viscosity $\nu$ that is characterized by a Reynolds number defined as $Re=UL/\nu$. Each dimension is uniformly discretized using $N$ cells leading to $\Delta x=L/(N-2)$ as two cells are reserved to represent solid boundaries on either side. As in Latt:2021:stlbm , we fix the time-step at $\Delta t=u_{LB}/U\Delta x$ where $u_{LB}$ is related to the Mach number ($Ma$) by the speed of sound in LBM units as $u_{LB}=Ma/c_{s}$ and is assumed here to be constant and fixed at $u_{LB}=0.06$ leading to $Ma\approx 0.1\ll 1$ supporting the incompressibility requirement. Simulations are run for $n$ iterations until a final dimensionless time $t=n(U/L)\Delta t$ is reached.
Figure 6 demonstrates the variation of horizontal ($u_{x}$) and vertical $(u_{z})$ components of the velocity field $u=(u_{x},u_{y},u_{z})$ at $Re=1000$ (steady flow), $Re=3200$ (unsteady flow) and $Re=10,000$ (turbulent flow). The velocity field is probed at the mid-plane associated with $y=0$ and normalized by the lid velocity $U$. Notice that the horizontal and vertical axes in this figure also represent the normalized (by $L/2$) distance along $x$ (for $u_{z}(x,0,0)/U$ on the y-axis) and along $z$ directions (for $u_{x}(0,0,z)/U$ on the x-axis) such that the clock-wise circulation of the flow inside the box can be visually apparent at this cross-section located at $y=0$. At $Re=3200$ and $Re=10000$, the presented results correspond with time-averaged values between $50\leq t\leq 250$ and $150\leq t\leq 500$ respectively similar to what has been reported in Latt:2021:stlbm . As shown in Figure 6 the results produced by XLB are in excellent agreement with all the reference data obtained from Albensoeder:2005:accurate ; Prasad:1989:reynolds ,
Figure 6: Comparing the variations of the velocity field against available reference data for 3D lid-driven cavity flow at (a) $Re=1000$, (b) $Re=3200$ and (c) $Re=10000$. In all panels the horizontal and vertical components of the velocity field ($u_{x},u_{z}$) are probed at mid-plane of the box (i.e. at $y=0$) and are normalized w.r.t. the lid velocity (note, $u_{lid}=(U,0,0)$). The reference data shown by colored circles are based on (a) spectral simulation results of Albensoeder:2005:accurate and (b, c) experimental results of Prasad:1989:reynolds . The BGK collision model together with the halfway bounce-back scheme was employed in (a, b) while the KBC model together with the regularized boundary scheme was necessarily chosen for (c) at $Re=10000$ to maintain numerical stability. In all cases $N=256$, and $u_{LB}=0.06$.
@@ 4.3 Flow over Cylinder
The geometric setup employed for this example is that reported in schafer1996benchmark (see their figure 1). Similar to lagrava2012advances , a Poiseuille profile with a mean speed of $U$ is imposed at the inlet such that $Re=100$ is achieved corresponding to a laminar but unsteady condition. We discretized the computational domain using $N=80$ cells across the cylinder diameter and fixed $\Delta t=u_{LB}/U\Delta x$ assuming $u_{LB}=0.003$ similar to lagrava2012advances .
@@
$C_{d}=\frac{2F_{d}}{\rho UD},\qquad C_{l}=\frac{2F_{d}}{\rho UD}$ (11)
where $F_{d}$ and $F_{l}$ are the drag and lift forces computed as surface integrals around the cylinder. Using the BGK model and the boundary conditions outlined earlier, the following maximum coefficients were obtained after $n$ iterations corresponding with dimensionless time $t=n(U/D)\Delta t=50$,
$C^{\text{max}}_{d}=3.25,\qquad C^{\text{max}}_{l}=0.930.$
These predicted values are sufficiently close to the reported range of $C_{d}^{\text{max}}=3.22-3.24$ and $C^{\text{max}}_{l}=0.99-1.01$ in schafer1996benchmark .
@@ 4.4 Turbulent Channel Flow
We now focus on a challenging and complex test associated with a continuously-forced wall-bounded turbulent flow inside a channel. This example aims to demonstrate XLB capabilities to perform accurate and efficient Large-Eddy Simulations (LES). Furthermore, this benchmark problem demonstrates how a body force can be added to the simulation using the exact-difference method of kupershtokh2004 .
@@
Figure 7 compares XLB results against Direct Numerical Simulation (DNS) results of Moser:1999:DNS at $Re_{\tau}=u_{\tau}h/\nu=180$. Here the distance is measured in wall units and defined as $y^{+}=yu_{\tau}/\nu$ while the averaged velocity (averaged along both periodic directions) is normalized using the wall friction velocity ($u_{\tau}=\tau_{w}/\rho$ where $\tau_{w}$ is the wall shear stress). For reference, we have also included the log-law of wall introduced by von Kármán as,
@@
Figure 7: (a) Comparing results taken with $h=50$ using XLB with DNS data of Moser:1999:DNS (solid line), as well as the log-law of wall (dashed line) (13). The KBC collision model Karlin:2014:gibbs in XLB was used for this example.
@@ 5 Performance
Compared to low-level programming languages like C++, it is not straightforward to incorporate certain performance optimizations in JAX without compromising the readability and flexibility of the codebase. In C++, for example, a user has direct control over many low-level aspects, such as memory layout, explicit SIMD (Single Instruction, Multiple Data) vectorization, and fine-grained control over how data is loaded into cache. These optimizations are required for achieving maximum performance in LBM. However, JAX abstracts away many of these low-level details to simplify the user experience and maintain its high-level API. While this abstraction makes JAX more user-friendly and flexible, especially for conducting complex mathematical operations and automatic differentiation, it can limit the ability to employ specific performance optimizations that are readily accessible to a C++ user.
To better understand the actual performance capabilities of XLB, especially in the context of these inherent trade-offs and JAX’s unique computational model, we have conducted a series of detailed performance evaluations. We discuss the performance characteristics of the XLB library across single GPU, multi-GPU, and distributed systems. The XLB library offers satisfactory performance even on desktop GPUs. Although it does not match the speed of highly optimized LBM codes written in low-level programming languages (see e.g. sailfish2014 ; meneghin2022neon ; krause2021openlb ; Latt:2021:stlbm ), it is sufficiently fast for most practical use cases. The library is also scalable, allowing for performance scaling when required.
All the performance evaluations are done based on a 3D lid-driven cavity simulation. The D3Q19 lattice and BGK collision model are used, and the simulation is executed without any I/O operations. Unless explicitly stated, all tests are done in single precision. We ignored the first iteration to account for the time required for the JIT compilation of the main loop. The performance is quantified in terms of million lattice updates per second (MLUPS) where a single lattice update corresponds with a full LBM iteration including collision, streaming and boundary condition steps.
The folder labeled examples/performance contains scripts used to generate the performance data presented in this section. Performance for any other simulation configuration can be evaluated by setting compute_MLUPS to True in the simulation configuration. It should also be noted that the library’s performance may change over time due to ongoing improvements or updates to dependencies like JAX or JAXLIB. Users are advised to test the library’s performance using the latest version.

### v3

@@ 3.6 LBM step
Leveraging JAX’s automatic differentiation capabilities, users can conveniently compute the gradient of a single LBM step with respect to its inputs. By applying the jax.grad function to the step method, one can directly obtain the gradients needed for sensitivity analysis, optimization, and machine learning – all without manual differentiation.
3.7 Mixed precision, distributed checkpointing, and I/O capabilities
The XLB library stands out with its support for mixed-precision computation. This allows users to tailor precision levels for computation and storage independently. For instance, a setting like f64/f32 indicates that the computation is performed in double precision (f64) while storage is handled in single precision (f32). Mixed-precision offers a balance of accuracy and efficiency: reduced memory usage enables larger simulations and potential performance gains, especially on modern GPUs due to their superior performance in lower precision arithmetic.
XLB also features robust checkpointing for long-running simulations. Users can save simulation states at intervals, facilitating recovery from system failures or allowing simulations to be paused and resumed at will. XLB’s distributed checkpointing system is optimized for simulations running across multiple devices or nodes.
The XLB library supports both Binary and ASCII VTK output formats through the PyVista library [77]. Integration with PhantomGaze [78] enables advanced, GPU-accelerated in-situ visualizations. This integration is streamlined by DLPack, which allows tensor data exchange between XLB and PhantomGaze directly on the GPU, eliminating traditional I/O bottlenecks. For a given simulation, users can customize the frequency of I/O and extract specific quantities of interest, such as velocity, pressure, lift, drag, energy spectrum, strain rate tensor, viscous dissipation and more.
We are actively expanding the capabilities of the XLB library. Please check out our repository for the latest updates and features.
@@ 4.1 Taylor-Green vortex in 2D
$u_{th}(\bm{x},t)$ $=$ $U_{0}\sin(k_{x}x)\cos(k_{y}y)\exp\left\{-\frac{t}{t_{d}}\right\}$
$v_{th}(\bm{x},t)$ $=$ $-U_{0}\cos(k_{x}x)\sin(k_{y}y)\exp\left\{-\frac{t}{t_{d}}\right\}$
$p_{th}(\bm{x},t)$ $=$ $1-\frac{\rho_{0}U_{0}^{2}}{4}\left[\cos(2k_{x}x)+\cos(2k_{y}y)\right]\exp\left\{-2\frac{t}{t_{d}}\right\}$ (9)
in which $k_{x}=2\pi/n_{x}$ and $k_{y}=2\pi/n_{y}$ with $n_{x}$ and $n_{y}$ indicating the number of cells in $x$ and $y$ directions, respectively. Also, $t_{d}$ denotes a diffusive timescale defined as $t_{d}=\left[2\nu(k_{x}^{2}+k_{y}^{2})\right]^{-1}$.
@@
Figure 5 illustrates $\epsilon_{u}$ (in panels a, c) and $\epsilon_{\rho}$ (in panels b, d) for a range of resolutions $N=\{32,\ldots,1024\}$ using either the BGK (in panels a, b) or the KBC (in panels c, d) models as the collision operator. The reported simulation results were conducted using various mixed precision ‘computation/storage’ pairs where f64 and f32 indicate double and single precision, respectively. For all cases, we used the D2Q9 lattice and periodic boundary conditions (recall that by default if no boundary condition is imposed, the streaming operation in XLB imposes periodic boundary conditions by construction). All the numerical results were obtained at $t_{f}=0.05t_{d}$ using a flow with $Re=1/\nu=1600$. We set $\Delta x=1/N$ and $\Delta t/\Delta x=0.04$ at $N=32$ where $\Delta t$ was decreased proportionally as $N$ was increased. The distribution functions $f_{i}$ were initialized based on the method of [79] using $\bm{u}_{th}(\bm{x},t=0)$ as per (9).
The expected second order of accuracy (i.e. $\epsilon_{u}\propto N^{-2}$ and $\epsilon_{\rho}\propto N^{-2}$) is obtained for both collision models using f64/f64 precision pair. It is worth mentioning that $\epsilon_{u}$ shows a more robust convergence based on the KBC model up to $N=1024$ while the convergence of $\epsilon_{u}$ degrades for the BGK model beyond $N=256$ for f64/f64. We may attribute this to the chosen time step size or the number of iterations used for the initialization process. If the storage precision is reduced to ‘f32’ (i.e. f64/f32 cases), the second order accuracy is preserved up to a lower resolution beyond which both $\epsilon_{u}$ and $\epsilon_{\rho}$ deteriorate as $N$ increases. This is due to the fact that the results become cluttered by numerical rounding errors and hence are not improved with increasing the resolution. This exact deterioration in accuracy occurs more vividly for even lower resolutions as the computational precision is also reduced to f32 (i.e. f32/f32 cases) leading to erroneous results for higher resolutions. Consequently, LBM results with f32/f32 are not reliable unless the distribution functions $f_{i}$ values are renormalized carefully to be centered around zero [80] (this feature is not currently available in XLB).
@@ 4.2 Lid-driven cavity in 3D
Here we investigate the 3D flow inside a lid-driven cavity to further verify XLB especially under steady-state, unsteady and turbulent scenarios with non-periodic boundaries. Namely, Dirichlet boundary conditions are employed to assign no-slip condition on all side walls except for the top lid which moves horizontally with a prescribed velocity of $\bm{u}(x,y,z=L/2)=(U,0,0)$. We consider a box of size $L^{3}$ containing an incompressible Newtonian fluid with viscosity $\nu$ that is characterized by a Reynolds number defined as $Re=UL/\nu$. Each dimension is uniformly discretized using $N$ cells leading to $\Delta x=L/(N-2)$ as two cells are reserved to represent solid boundaries on either side. As in [81], we fix the time-step at $\Delta t=u_{LB}/U\Delta x$ where $u_{LB}$ is related to the Mach number ($Ma$) by the speed of sound in LBM units as $u_{LB}=Ma/c_{s}$ and is assumed here to be constant and fixed at $u_{LB}=0.06$ leading to $Ma\approx 0.1\ll 1$ supporting the incompressibility requirement. Simulations are run for $n$ iterations until a final dimensionless time $t=n(U/L)\Delta t$ is reached.
Figure 6 demonstrates the variation of horizontal ($u_{x}$) and vertical $(u_{z})$ components of the velocity field $\bm{u}=(u_{x},u_{y},u_{z})$ at $Re=1000$ (steady flow), $Re=3200$ (unsteady flow) and $Re=10,000$ (turbulent flow). The velocity field is probed at the mid-plane associated with $y=0$ and normalized by the lid velocity $U$. Notice that the horizontal and vertical axes in this figure also represent the normalized (by $L/2$) distance along $x$ (for $u_{z}(x,0,0)/U$ on the y-axis) and along $z$ directions (for $u_{x}(0,0,z)/U$ on the x-axis) such that the clock-wise circulation of the flow inside the box can be visually apparent at this cross-section located at $y=0$. At $Re=3200$ and $Re=10000$, the presented results correspond with time-averaged values between $50\leq t\leq 250$ and $150\leq t\leq 500$ respectively similar to what has been reported in [81]. As shown in Figure 6 the results produced by XLB are in excellent agreement with all the reference data obtained from [82, 83],
Figure 6: Comparing the variations of the velocity field against available reference data for 3D lid-driven cavity flow at (a) $Re=1000$, (b) $Re=3200$ and (c) $Re=10000$. In all panels the horizontal and vertical components of the velocity field ($u_{x},u_{z}$) are probed at mid-plane of the box (i.e. at $y=0$) and are normalized w.r.t. the lid velocity (note, $u_{lid}=(U,0,0)$). The reference data shown by colored circles are based on (a) spectral simulation results of [82] and (b, c) experimental results of [83]. The BGK collision model together with the half-way bounce-back scheme was employed in (a, b) while the KBC model together with the regularized boundary scheme was necessarily chosen for (c) at $Re=10000$ to maintain numerical stability. In all cases $N=256$, and $u_{LB}=0.06$.
@@ 4.3 Flow over cylinder
  • 1.
    Lift and drag computations using the momentum exchange method of [84] as described in [85],
  • 2.
    The no-slip boundary condition for curved geometries based on [86],
  • 3.
    Inflow boundary condition using either the Zou-He [75] or the regularized approach [76],
  • 4.
    Outflow boundary condition based on the extrapolation scheme of [69].
Figure 7: Illustrating the geometric setup for the flow over 2d cylinder benchmark example.
The geometric setup employed for this example is shown in Figure 7 and is identical to that reported in [87] (see their figure 1). Similar to [88], a Poiseuille profile with a mean speed of $U$ is imposed at the inlet such that $Re=100$ is achieved corresponding to a laminar but unsteady condition. We discretized the computational domain using $N=80$ cells across the cylinder diameter and fixed $\Delta t=u_{LB}/U\Delta x$ assuming $u_{LB}=0.003$ similar to [88].
@@
$C_{d}=\frac{2F_{d}}{\rho U^{2}D},\qquad C_{l}=\frac{2F_{l}}{\rho U^{2}D}$ (11)
where $F_{d}$ and $F_{l}$ are the drag and lift forces computed as surface integrals around the cylinder. Using the BGK model and the boundary conditions outlined earlier, the following maximum coefficients were obtained after $n$ iterations corresponding with dimensionless time $t=n(U/D)\Delta t=100$,
$C^{\text{max}}_{d}=3.25,\qquad C^{\text{max}}_{l}=0.985.$
Furthermore, the vortex shedding in this example may be characterized by the Strouhal number defined as $St=f_{s}D/U$ in which $f_{s}$ is the dominant shedding frequency. Using the time series associated with $C_{l}(t)$ for $50\leq t\leq 100$, we obtained $St=0.300$. These predicted values are in good agreement with the reported range of $C_{d}^{\text{max}}=3.22-3.24$, $C^{\text{max}}_{l}=0.99-1.01$ and $St=0.295-0.305$ in [87] which are based on a large collection of simulations using various numerical techniques.
@@ 4.4 Turbulent channel flow
We now focus on a challenging and complex test associated with a continuously-forced wall-bounded turbulent flow inside a channel. This example aims to demonstrate XLB capabilities to perform accurate and efficient Large-Eddy Simulations (LES). Furthermore, this benchmark problem demonstrates how a body force can be added to the simulation using the exact-difference method of [89].
@@
Figure 8 compares XLB results against Direct Numerical Simulation (DNS) results of [90] at $Re_{\tau}=u_{\tau}h/\nu=180$. Here the distance is measured in wall units and defined as $y^{+}=yu_{\tau}/\nu$ while the averaged velocity (averaged along both periodic directions) is normalized using the wall friction velocity ($u_{\tau}=\tau_{w}/\rho$ where $\tau_{w}$ is the wall shear stress). For reference, we have also included the log-law of wall introduced by von Kármán as,
@@
Figure 8: (a) Comparing results taken with $h=50$ using XLB with DNS data of [90] (solid line), as well as the log-law of wall (dashed line) (13). The KBC collision model [71] in XLB was used for this example.
@@ 5 Performance
Compared to low-level programming languages like C++, it is not straightforward to incorporate certain performance optimizations in JAX without compromising the readability and flexibility of the codebase. In C++, for example, a user has direct control over many low-level aspects, such as memory layout, explicit SIMD (Single Instruction, Multiple Data) vectorization, and fine-grained control over how data is loaded into cache. These optimizations are required for achieving maximum performance in LBM. However, JAX abstracts away many of these low-level details to simplify the user experience and maintain its high-level API. While this abstraction makes JAX more user-friendly and flexible, especially for conducting complex mathematical operations and automatic differentiation, it can limit the ability to employ specific performance optimizations that are readily accessible to a C++ user. We are actively working to mitigate these limitations in XLB and achieve state-of-the-art performance and aim to add a kernel-based backend to XLB based on Warp [91]. Below, we present performance characteristics of the current XLB version, which relies on the JAX backend.
To better understand the actual performance capabilities of XLB, especially in the context of these inherent trade-offs and JAX’s unique computational model, we have conducted a series of detailed performance evaluations. We discuss the performance characteristics of the XLB library across single GPU, multi-GPU, and distributed systems. The XLB library offers satisfactory performance even on desktop GPUs. Although it does not match the speed of highly optimized LBM codes written in low-level programming languages (see e.g. [92, 67, 66, 81]), it is sufficiently fast for most practical use cases. The library is also scalable, allowing for performance scaling when required.
All the performance evaluations are done based on a 3D lid-driven cavity simulation with stationary walls on all sides except the top lid that moves horizontally at a constant velocity. The full-way bounce-back scheme was used to impose $\bm{u}=(0,0,0)$ and the equilibrium boundary condition was employed for the top lid to impose $\bm{u}=(U,0,0)$. The D3Q19 lattice and BGK collision model are used, and the simulation is executed without any I/O operations. Unless explicitly stated, all tests are done in single precision. We ignored the first iteration to account for the time required for the JIT compilation of the main loop. The performance is quantified in terms of million lattice updates per second (MLUPS) where a single lattice update corresponds with a full LBM iteration including collision, streaming and boundary condition steps.
The examples folder in the repository contains scripts that can be used to reproduce the performance data presented in this section. It should also be noted that the library’s performance may change over time due to ongoing improvements or updates to dependencies like JAX or JAXLIB. Users are advised to test the library’s performance using the latest version.